OpenStudy (mikey):

will medal for answer Simplify (6 + i)(8 − 3i). 51 − 10i 48 − 3i2 45 − 10i 48 − 10i − 3i2

OpenStudy (mikey):

I know its between b and d I'm pretty sure

OpenStudy (mikey):

@3mar

OpenStudy (3mar):

It is the first choice. Let me show you

OpenStudy (3mar):

(6+i)(8-3i) =48-18i+8i-3i^2 =48-10i+3 =51-10i Hope that helps

OpenStudy (3mar):

@mikey

OpenStudy (mikey):

howd the 3i turn into a 3

OpenStudy (mikey):

@3mar

OpenStudy (mikey):

because u squared it?

OpenStudy (3mar):

\[(6+i)(8-3i)=48-18i+8i-3i^2=48-10i-3i^2\] and of course you know that \[i^2=-1\] because \[i=\sqrt{-1}\] Good?

OpenStudy (3mar):

so\[3i^2=3(-1)=-3\] and proceed

OpenStudy (mikey):

good explanation

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (mikey):

np

OpenStudy (3mar):

Good learner! I really like who ask and stop the explanation when he sticks to something, even it looks like so simple